\begin{thebibliography}{99}
%\bibitem[AG62]{AG62}	A. Andreotti and H. Grauert, \emph{Th\'eor\`eme de finitude pour la cohomologie des espaces complexes}, Bull. Soc. Math. France \textbf{90}, (1962), 193-259.
 %\bibitem[AV65]{AV65} A.~Andreotti and E.~Vesentini, \emph{Carleman estimates for the Laplace-Beltrami equation in complex manifolds}, Publ. Math. I.H.E.S. \textbf{25}, (1965), 81-130.
 %\bibitem[Ber98]{Ber98} B.~Berndtsson, \emph{Prekopa's theorem and Kiselman's minimum principle for plurisubharmonic functions}, Math. Ann. \textbf{312} (1998), 785--792. 
% \bibitem[Ber06]{Ber06} B.~Berndtsson, \emph{Subharmonicity properties of the Bergman kernel and some other functions associated to pseudoconvex domains}, Ann. Inst. Fourier, \textbf{56} (2006), 1633--1662.
% \bibitem[Ber09]{Ber09} B.~Berndtsson, \emph{Curvature of vector bundles associated to holomorphic fibrations}, Ann. of Math. \textbf{169} (2009), no.~2, 531--560. 
 %\bibitem[Ber10]{Ber10} B.~Berndtsson, \emph{An introduction to things $\dbar$, Analytic and algebraic geometry},IAS/Park City Math. Ser., vol. 17, Amer. Math. Soc., Providence, RI, 2010, pp. 7-76.
 %\bibitem[Ber11]{Ber11} B.~Berndtsson, \emph{Strict and nonstrict positivity of direct image bundles}, Math. Z. \textbf{269} (2011), 1201--1218.
 %\bibitem{Ber20} B.~Berndtsson, \emph{Lelong numbers and vector bundles}, J. Geom. Anal. \textbf{30} (2020), 2361--2376. 
 %\bibitem{BP08} B.~Berndtsson and M.~P{\u{a}}un, \emph{Bergman kernels and the pseudoeffectivity of relative canonical bundles}, Duke Math. J. \textbf{145} (2008), no.~2, 341--378.
 %\bibitem[BP10]{BP} B.~Berndtsson and M.~P\u{a}un, \emph{Bergman kernels and subadjunction}, arXiv:1002.4145.
% \bibitem[B\l o13]{Blo13} Z.~B\l ocki, \emph{Suita conjecture and the Ohsawa-Takegoshi extension theorem}, Invent. Math. \textbf{193}, (2013), 149--158. 
 %\bibitem[BL76]{BL76} H.~J.~Brascamp and E.~H.~Lieb, \emph{On extensions of the Brunn-Minkowski and Pr\'ekopa-Leindler theorems, including inequalities for log concave functions, and with an application to the diffusion equation}, J. Funct. Anal. \textbf{22}, (1976), no.~4, 366-389. 
 %\bibitem[CF90]{CF90} F. Campana and H. Flenner, \emph{A characterization of ample vector bundles on a curve}, Math. Ann. \textbf{287}, (1990), no.~4, 571-575.
 %\bibitem[CP17]{CP17} J. Cao and M.~P\u{a}un, \emph{Kodaira dimension of algebraic fiber spaces over abelian varieties}, Invent. Math. \textbf{207}, (2017), 345-387.
%\bibitem[Cor05]{Cor05} D. Cordero-Erausquin, \emph{On Berndtsson's generalization of Pr\'ekopa's theorem}, Math. Z. \textbf{249}, (2005), 401-410.
 %\bibitem{deC98} M.~A.~A. de~Cataldo, \emph{Singular {H}ermitian metrics on vector bundles}, J. Reine Angew. Math. \textbf{502} (1998), 93--122.
 %\bibitem{Dem82} J.-P. Demailly, \emph{Estimations $L^2$ pour l'op\'erateur $\dbar$ d'un fibr\'e vectoriel holomorphe semi-positif au dessus d'une vari\'et\'e K\"ahl\'erienne compl\`ete}, Ann. Sci. Ec. Norm. Sup. \textbf{15} (1982), 457--511.
 %\bibitem[Dem93]{Dem93} J.-P. Demailly, \emph{A numerical criterion for very ample line bundles}, J. Differential Geom. \textbf{37}, (1993), 323-374. 
 %\bibitem[Dem]{Dem12} J.-P. Demailly, \emph{Analytic methods in algebraic geometry}, Surveys of Modern Mathematics, vol.~1, International Press, Somerville, MA; Higher Education Press, Beijing, 2012.
 %\bibitem[Dem-book]{Dem-agbook} J.-P. Demailly, \emph{Complex analytic and differential geometry}, \url{http://www-fourier.ujf-grenoble.fr/~demailly/manuscripts/agbook.pdf}.
%\bibitem{DS} J.-P. Demailly and H.~Skoda, \emph{Relations entre les notions de positivit\'e de P.~A.~Griffiths et de S.~Nakano}, S\'eminaire P.~Lelong-H.~Skoda (Analyse), ann\'ee 1978/79, Lecture notes in Math. no \textbf{822}, Springer-Verlag, Berlin (1980) 304--309.
 \bibitem{DNW21} F.~Deng, J.~Ning, and Z.~Wang, \emph{Characterizations of plurisubharmonic functions}, Sci. China Math. \textbf{64} (2021), 1959--1970.
 \bibitem{DNWZ23} F.~Deng, J.~Ning, Z.~Wang, and X.~Zhou, \emph{Positivity of holomorphic vector bundles in terms of $L^p$-estimates for $\overline{\partial}$}, Math. Ann. \textbf{385} (2023), 575--607. %https://doi.org/10.1007/s00208-021-02348-7.
 %\bibitem[DWZZ18]{DWZZ18} F.~Deng, Z.~Wang, L.~Zhang, and X.~Zhou, \emph{New characterizations of plurisubharmonic functions and positivity of direct image sheaves}, arXiv:1809.10371.
 %\bibitem[DWZZ19]{DWZZ19} F.~Deng, Z.~Wang, L.~Zhang, and X.~Zhou, \emph{A new proof of Kiselman's minimum principle for plurisubharmonic functions}, Comptes Rendus Math\'ematique \textbf{357}, (2019), no.~4, 345-348. 
 %\bibitem[DWZZ20]{DWZZ20} F.~Deng, Z.~Wang, L.~Zhang, and X.~Zhou, \emph{Linear invariants of complex manifolds and their plurisubharmonic variations}, J. Funct. Anal. \textbf{279}, (2020), https://doi.org/10.1016/j.jfa.2020.108514. 
 %\bibitem[DZZ14]{DZZ14} F.~Deng, H.~Zhang, and X.~Zhou, \emph{Positivity of direct images of positively curved volume forms}, Math. Z. \textbf{278} (2014), no.~1-2, 347--362. 
 %\bibitem[Gri69]{Gri69}P. A. Griffiths, \emph{Hermitian differential geometry, Chern classes, and positive vector bundles}, Global Analysis, papers in honor of K. Kodaira, Princeton Univ. Press, Princeton, (1969), 181-251.
\bibitem{GZ15} Q.~Guan and X.~Zhou, \emph{A solution of an $L^2$ extension problem with an optimal estimate and applications}, Ann. of Math. (2) \textbf{181} (2015), no.~3, 1139--1208.
 \bibitem{HPS18} C.~Hacon, M.~Popa, and C.~Schnell, \emph{Algebraic fiber spaces over Abelian varieties: around a recent theorem by Cao and Paun}, Contemp. Math. \textbf{712} (2018), 143--195. 
 \bibitem{Hor65} L.~H\"ormander, \emph{$L^2$ estimates and existence theorems for the $\dbar$ operator}, Acta Math. \textbf{113} (1965), 89--152. 
 %\bibitem{Hos17} G.~Hosono, \emph{Approximations and examples of singular Hermitian metrics on vector bundles}, Ark. Mat. \textbf{55} (2017), 131--153.
% \bibitem[Hos19]{Hos19} G.~Hosono, \emph{On sharper estimates of Ohsawa--Takegoshi $L^2$-extension theorem}, J. Math. Soc. Japan \textbf{71} (2019), no.~3, 909--914.
 %\bibitem[HI21]{HI21} G.~Hosono and T.~Inayama, \emph{A converse of H\"ormander's $L^2$-estimate and new positivity notions for vector bundles}, Sci. China Math. \textbf{64} (2021), 1745--1756. %https://doi.org/10.1007/s11425-019-1654-9.
 %\bibitem{InaJGA} T.~Inayama, \emph{Curvature currents and Chern forms of singular Hermitian metrics on holomorphic vector bundles}, J. Geom. Anal. \textbf{30} (2020), 910--935. 
 %\bibitem{Ina18} T.~Inayama, \emph{$L^2$ estimates and vanishing theorems for holomorphic vector bundles equipped with singular Hermitian metrics}, Michigan Math. J. \textbf{69} (2020), 79--96.
 %\bibitem{Ina20} T.~Inayama, \emph{Nakano positivity of singular Hermitian metrics and vanishing theorems of Demailly-Nadel-Nakano type}, Algebraic Geometry \textbf{9} (2022), 69--92. 
\bibitem{Ina22a} T.~Inayama, \emph{Optimal $L^2$-extensions on tube domains and a simple proof of Pr\'ekopa's theorem}, J. Geom. Anal. \textbf{32}, 32 (2022). \url{https://doi.org/10.1007/s12220-021-00796-w}.
 \bibitem{Ina22b} T.~Inayama, \emph{$L^2$-extension indices, sharper estimates and curvature positivity}, arXiv:2210.08456. 
 %\bibitem{Iwa21} M.~Iwai, \emph{Nadel-Nakano vanishing theorems of vector bundles with singular Hermitian metrics},  Annales de la Facult\'e des sciences de Toulouse: Math\'ematiques (6), \textbf{30} (2021), no.~1, 63--81.
% \bibitem[KP21]{KP21} A.~Khare and V.~Pingali, \emph{On an asymptotic characterisation of Griffiths semipositivity}, Bull. des Sci. Math. \textbf{167} (2021), 102956.
% \bibitem[Kik22]{Kik22} S.~Kikuchi, \emph{On sharper estimates of Ohsawa--Takegoshi $L^2$-extension theorem in higher dimensional case}, manuscripta math. (2022). https://doi.org/10.1007/s00229-021-01366-8.
% \bibitem[Kis78]{Kis78} C.~Kiselman, \emph{The partial Legendre transformation for plurisubharmonic functions}, Invent. Math. \textbf{49} (1978), no.~2, 137--148. 
 %\bibitem{LRRS18} R.~L\"ark\"ang, H.~Raufi, J.~Ruppenthal and M.~Sera, \emph{Chern forms of singular metrics on vector bundles}, Adv. Math. \textbf{326} (2018), 465--489. 
 %\bibitem{LSY13} K.-F.~Liu, X.-F. Sun and X.-K. Yang, \emph{Positivity and vanishing theorems for ample vector bundles}, J. Algebraic Geom. \textbf{22} (2013), 303--331.   
 \bibitem{LZ23} Z.~Liu and W.~Xu, \emph{Characterizations of Griffiths positivity, pluriharmonicity and flatness}, preprint.
 %\bibitem[LY14]{LY14} K. Liu and X. Yang, \emph{Curvatures of direct image sheaves of vector bundles and applications}, J. Differential Geom. \textbf{98}, (2014), 117-145.
 %\bibitem[MT08]{MT08} C. Mourougane and S. Takayama, \emph{Hodge metrics and the curvature of higher direct images}, Ann. Sci. \'Ec. Norm. Sup\'er. \textbf{41}, (2008), 905-924.
 %\bibitem[MT09]{MT09} C. Mourougane and S. Takayama, \emph{Extension of twisted Hodge metrics for K\"ahler morphisms}, J. Differential Geom. \textbf{83}, (2009) 131-161.
 %\bibitem{Nad90} A.~M.~Nadel, \emph{Multiplier ideal sheaves and K\"ahler-Einstein metrics of positive scalar curvature}, Ann. of Math. (2) \textbf{132} (1990), no.~3, 549--596.
 %\bibitem[NZZ16]{NZZ} J.~Ning, H.~Zhang, and X.~Zhou, \emph{On $p$-Bergman kernel for bounded domains in $\mathbb{C}^n$}, Comm. Anal. Geom. \textbf{24}, (2016), no.~4, 887-900.
 \bibitem{OT87} T.~Ohsawa and K.~Takegoshi, \emph{On the extension of $L^2$ holomorphic functions}, Math. Z. \textbf{195} (1987), no.~2, 197--204. 
 %\bibitem{PT18} M.~P{\u{a}}un and S.~Takayama, \emph{Positivity of twisted relative pluricanonical bundles and their direct images}, J. Algebraic Geom. \textbf{27} (2018), 211--272. 
 %\bibitem[Pr\'e73]{Pre73} A.~Pr\'ekopa, \emph{On logarithmic concave measures and functions}, Acta. Sci. Math. (Szeged) \textbf{34}, (1973), 335-343. 
 %\bibitem{Rau15} H.~{Raufi}, \emph{Singular hermitian metrics on holomorphic vector bundles}, Ark. Mat. \textbf{53} (2015), no.~2, 359--382.
%\bibitem{Siu76} Y.~T.~Siu, \emph{Every Stein subvariety admits a Stein neighborhood}, Invent. Math. \textbf{38} (1976/77), 89--100.
%\bibitem[Siu98]{Siu98} Y.~T.~Siu, \emph{Invariance of plurigenera}, Invent. Math. \textbf{134} (1998), 661--673.
%\bibitem[Sko72]{Sko72}H.~Skoda, \emph{Sous-ensembles analytiques d'ordre fini ou infini dans $\mathbb{C}^n$}, Bull. Soc. Math. France, \textbf{100} (1972), 353-408.
%\bibitem[Ume73]{Ume73} H. Umemura, \emph{Some results in the theory of vector bundles}, Nagoya Math. J. \textbf{52}, (1973), 97-128.
%\bibitem[Xu22]{Xu22} W.~Xu, \emph{A Quantitative Characterization of Griffiths Positivity}, arXiv:2210.17361. 
%\bibitem[XZ22]{XZ22} W.~Xu and Z.~Zhou, \emph{Optimal $L^2$ Extensions of Openness Type}, arXiv:2202.04791.
%\bibitem[Yan18]{Yan18} X. Yang, \emph{RC-positivity, rational connectedness and Yau's conjecture}, Camb. J. Math. \textbf{6}, (2018), no.~2, 183-212. 
%\bibitem[Yan19]{Yan19} X. Yang, \emph{A partial converse to the Andreotti-Grauert theorem}, Compos. Math. \textbf{155}, (2019), 89-99.
\end{thebibliography}

\documentclass[leqno,12pt]{amsart} 
\usepackage
[top=30truemm,bottom=30truemm,left=25truemm,right=25truemm]
{geometry}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amscd}
\usepackage{mathrsfs}
\usepackage{graphicx}
\usepackage{url}
  \makeatletter

\@addtoreset{equation}{section}
\@namedef{subjclassname@2020}{\textup{2020} Mathematics Subject Classification}
\makeatother
\theoremstyle{plain} 
\newtheorem{theorem}{\indent\bf Theorem}[section]

\newtheorem{proposition}[theorem]{\indent\bf Proposition}

\newtheorem{conjecture}[theorem]{\indent\bf Conjecture}
\theoremstyle{definition} 
\newtheorem{definition}[theorem]{\indent\bf Definition}

\newcommand{\dbar}{\overline{\partial}}

\newcommand{\R}{\mathbb{R}}

\newcommand{\C}{\mathbb{C}}

\newcommand{\B}{\mathbb{B}}

\newcommand{\Rea}{\mathrm{Re}}

\begin{document}
\pagestyle{plain}
\thispagestyle{plain}

\title[Note on characterizing pluriharmonic functions]
{A note on characterizing pluriharmonic functions via the Ohsawa--Takegoshi extension theorem}

\author[T. INAYAMA]{Takahiro INAYAMA}
\address{Department of Mathematics\\
Faculty of Science and Technology\\
Tokyo University of Science\\
2641 Yamazaki, Noda\\
Chiba, 278-8510\\
Japan
}
\email{inayama\_takahiro@rs.tus.ac.jp}
\email{inayama570@gmail.com}
\subjclass[2020]{32A10, 32A36}
\keywords{ 
Ohsawa--Takegoshi extension theorem, $L^2$-extension, plurisubharmonic function, pluriharmonic function, $L^2$-extension index.
}

\begin{abstract}
 For a continuous function, we prove that the function is pluriharmonic if and only if the equality part of the optimal Ohsawa--Takegoshi $L^2$-extension theorem is satisfied with respect to the metric having the function as a weight.  
 This partially resolves the conjecture proposed by the author. 
\end{abstract}


\maketitle
\setcounter{tocdepth}{2}

\section{Introduction}\label{sec:intro}

On a one-dimensional complex domain, as is well known, a subharmonic function is characterized by the mean value inequality, and when the equality of the inequality holds, the function becomes a harmonic function. Subharmonic functions play an important role in complex analysis and geometry, and are also used in important theorems such as H\"ormander's $L^2$-estimate \cite{Hor65} and Ohsawa--Takegoshi's $L^2$-extension theorem \cite{OT87}.

On the other hand, recent research has revealed that the fact that the optimal Ohsawa--Takegoshi $L^2$-extension theorem holds itself guarantees the subharmonicity of the weight. This property is called the minimal extension property \cite{HPS18} or the optimal $L^p$-extension property \cite{DNW21,DNWZ23} in a general setting, and has been widely studied and applied by various experts (cf. \cite{GZ15,HPS18,DNW21,DNWZ23,Ina22a,Ina22b}). In other words, subharmonic functions can be characterized by the inequality part of the optimal Ohsawa--Takegoshi $L^2$-extension theorem. 
Based on the analogy with the above, we propose the following conjecture in \cite[Appendix A]{Ina22b}: 

\begin{conjecture}\label{conj:pluriharmonic}
Let $\varphi$ be an upper semi-continuous function on a domain $\Omega\subset \C^n$. 
Then the following are equivalent: 
\begin{enumerate}
    \item $\varphi$ is pluriharmonic.
    \item $\varphi>-\infty$ and for any holomorphic cylinder $a+P_{r,s,A}$, where $(a,r,s,A)\in \Omega_{\widetilde{\delta}}$ {\rm (}see the notation below{\rm )}, there exists a unique holomorphic function $f$ on $a+P_{r,s,A}$ satisfying $f(a)=1$ and 
    \[
    \int_{a+P_{r,s,A}} |f|^2e^{-\varphi} \leq |P_{r,s,A}|e^{-\varphi(a)},
    \]
    where $|P_{r,s,A}|$ is the volume of $P_{r,s,A}$. 
\end{enumerate}
\end{conjecture}


We proved this conjecture in the case that $\varphi$ was smooth by using the theory of the $L^2$-extension index in \cite{Ina22b}. 
Our aim of this paper is to give an affirmative answer to the above conjecture for a \textit{continuous} function $\varphi$. 
To clarify the claim of the theorem, several symbols will be prepared. 
Let $\Delta_r=\{ z\in \C \mid |z|<r\}$, $\B^m_s=\{ z\in\C^m\mid |z|<s \}$ and $P_{r,s,A}=A(\Delta_r\times \B^{n-1}_s)$ for $r,s>0$ and $A\in \mathbf{U}(n)$, where $\mathbf{U}(n)$ is the set of all $n$-dimensional unitary groups.
Here $P_{r,s,A}$ is a holomorphic cylinder.
We let $\Omega_{\widetilde{\delta}}=\{ (a,r,s,A)\in \Omega\times\R_{>0}\times\R_{>0}\times \mathbf{U}(n)\mid a+P_{r,s,A}\subset \Omega \}$. 
Then the main theorem can be stated as follows. 

\begin{theorem}\label{mainthm:pluriharmonic}
Let $\varphi$ be a \textbf{continuous} function on $\Omega$. Then $\varphi$ is pluriharmonic if and only if for any $(a,r,s,A)\in \Omega_{\widetilde{\delta}}$, there exists a \textit{unique} holomorphic function $f$ on $a+P_{r,s,A}$ satisfying $f(a)=1$ and 
\[
\int_{a+P_{r,s,A}} |f|^2e^{-\varphi}\leq |P_{r,s,A}| e^{-\varphi(a)}.
\]
\end{theorem}

Note that the above $f$ satisfies the equality 
\[
\int_{a+P_{r,s,A}} |f|^2e^{-\varphi}= |P_{r,s,A}| e^{-\varphi(a)}.
\]
For the proof, we use the terms of the $L^2$-extension index and the characterization of log-plurisubharmonicity.



\vskip3mm
{\bf Acknowledgment. } 
The author is supported by Japan Society for the Promotion of Science, Grant-in-Aid for Research Activity Start-up (Grant No. 21K20336) and Grant-in-Aid for Early-Career Scientists (Grant No. 23K12978).
He also expresses his gratitude to Dr. Wang Xu for fruitful discussions, sharing various ideas and his generous response.

\section{The proof of main theorem}

In order to prove the main result, we prepare several notions. 
First, we review the minimal extension property \cite{HPS18} or the optimal $L^2$-extension property \cite{DNW21,DNWZ23}.
In this paper, we follow the formulation of Deng, Ning, Wang and Zhou.

\begin{definition}[{\cite[Definition 1.1]{DNW21}}]\label{def:optimall2}
Let $\Omega$ be a domain in $\C^n$ and $\varphi$ be an upper semi-continuous function. 
Then we say that $\varphi$ satisfies {\it the optimal $L^2$-extension property} if for any $(a,r,s,A)\in \Omega_{\widetilde{\delta}}$, there exists a holomorphic function $f$ on $a+P_{r,s,A}$ such that $f(a)=1$ and 
\[
\int_{a+P_{r,s,A}} |f|^2e^{-\varphi}\leq |P_{r,s,A}|e^{-\varphi(a)}.
\]
\end{definition}

The following theorem is important in relation to this definition.

\begin{theorem}[{\cite[Theorem 1.6]{DNW21}}]\label{thm:optimall2}
    Keep the setting. 
    Then $\varphi$ is plurisubharmonic if and only if $\varphi$ satisfies the optimal $L^2$-extension property. 
\end{theorem}

Next, we explain the $L^2$-extension index introduced by the author in \cite{Ina22b}. 
Here we adopt a slightly extended definition.

\begin{definition}\label{def:L2indexgeneral}
	Let $\varphi$ be a function $\varphi\colon\Omega\to [-\infty, \infty]$.
	Then we define the \textit{$L^2$-extension index} $L_\varphi$ of $\varphi$ on $\Omega_{\widetilde{\delta}}$ as follows: for $(a,r,s,A)\in \Omega_{\widetilde{\delta}}$, if $-\infty <\varphi(a)<+\infty$,
	\begin{align*}
	L_{\varphi}(a,r,s,A):&= \frac{1}{|P_{r,s,A}|K_{P_{r,s,A}, \varphi}(a)}\\
	&=\inf \left\{ \frac{\int_{a+P_{r,s,A}}|f|^2e^{-\varphi}}{|P_{r,s,A}|e^{-\varphi(a)}} ~\middle|~ f\in A^2(a+P_{r,s,A}, \varphi) \And f(a)=1 \right\},
	\end{align*}
	if $\varphi(a)=+\infty$, $L_\varphi(a,r,s,A)=+\infty$ and if $\varphi(a)=-\infty$, $L_{\varphi}(a,r,s,A)=0$.
	Here, $K_{{P_{r,s,A}},\varphi}$ is the weighted Bergman kernel on $P_{r,s,A}$ with respect to $\varphi$, $A^2(a+P_{r,s,A}, \varphi)=\{ f\in\mathcal{O}(a+P_{r,s,A})\mid \int_{a+P_{r,s,A}} |f|^2e^{-\varphi}<+\infty \}$ and $\mathcal{O}(a+P_{r,s,A})$ is the space of all holomorphic functions on $a+P_{r,s,A}$.
\end{definition}

By using this notion, we can rephrase the optimal $L^2$-extension property as follows: if $L_\varphi \leq 1$ for an upper semi-continuous function, then $\varphi$ is plurisubharmonic. 
Our goal is to prove the next proposition that is paired with the above result. 

\begin{proposition}\label{prop:lsc}
    Let $\varphi$ be a lower semi-continuous function on a domain $\Omega\subset \C^n$ with $\varphi \not\equiv +\infty$. 
    If $L_\varphi\geq 1$, then $\varphi$ is plurisuperharmonic.
\end{proposition}



\begin{proof}
    We use the proposition that for a non-negative function $v$ with $v\not \equiv 0$, $\log v$ is plurisubharmonic if and only if $v e^{2\Rea g}$ is plurisubharmonic for every polynomial $g$. 
    We take any polynomial $g$ and any $(a,r,s,A)\in \Omega_{\widetilde{\delta}}$. We may assume that $\varphi(a)<+\infty$. 
    If $L_\varphi \geq 1$, it holds that
    \[
    \int_{a+P_{r,s,A}}|e^g|^2e^{-\varphi}\geq |P_{r,s,A}| |e^{g(a)}|^2e^{-\varphi(a)}.
    \]
    Since $-\varphi$ is upper semi-continuous, we can say that $e^{-\varphi}e^{2\Rea g}$ is plurisubharmonic (see Lemma 3.1 in \cite{DNW21}). Hence, $-\varphi$ is plurisubharmonic. 
\end{proof}

By using this proposition, we can prove Theorem \ref{mainthm:pluriharmonic}. 

\begin{proof}[Proof of Theorem \ref{mainthm:pluriharmonic}]
If $\varphi$ is pluriharmonic, take a holomorphic function $h$ on $a+P_{r,s,A}$ satisfying $2\Rea (h)=\varphi$ and use the argument in \cite[Section 5]{Ina22b}. 
Then we only prove the if part. 
Note that the assumption in Theorem \ref{mainthm:pluriharmonic} says that $L_\varphi\equiv 1$. 
	Since $L_\varphi\leq 1$ and $\varphi$ is upper semi-continuous, Theorem \ref{thm:optimall2} implies $\varphi$ is plurisubharmonic. 
	Also, since $L_\varphi\geq 1$ and $\varphi$ is lower semi-continuous, Proposition \ref{prop:lsc} implies that $\varphi$ is plurisuperharmonic.
\end{proof}

I have discussed Conjecture \ref{conj:pluriharmonic} with Wang Xu. 
On May 15, I sent him the above proof. 
Then, by using the linearity of certain minimal $L^2$-integrals, Xu showed that the assumption that $\varphi$ is lower semi-continuous is not needed and sent me the proof on May 16. 
About a month later, with Zhuo Liu, Xu consequently obtained the result that the upper semi-continuity of $\varphi$ is also unnecessary and sent the manuscript \cite{LZ23} to me on June 26. 
Although their result is literally a generalization of my theorem, they encouraged me to write this paper on my result as a first important step.
I would like to thank them for their consideration and warm encouragement.



\end{document}